Determine the ratio of the frequency (use scientific notation and 3 significant figures) between visible lights at a wavelength of 400 nm against visible light at a wavelength of 750 nm.
Answer #1

From the concept of electromagnetic radiation, C = \nu \lambda

c = light velocity

\nu = frequency of photon

\lambda = wave length of photon

For two different lights of different wave-lengths and frequencies,

\nu1*\lambda1 = \nu 2*\lambda2

therfore, \nu 1/\nu2 = \lambda 2/\lambda1 = 750/400 = 1.875 = 188*10^2